# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ Abstract
Chebyshev Filtered Subspace Iteration (ChFSI) has emerged as a robust alternative to Krylov eigensolvers for extracting a small subset of extremal eigenpairs from large sparse matrices, particularly in situations where these eigenpairs must be computed repeatedly as the system matrix evolves within an outer iteration. Such scenarios arise prominently in computational physics, including ab initio materials modelling using density functional theory (DFT), as well as in certain stability and coupled multiphysics analyses where eigenvalue information augments nonlinear solution procedures. In this work, we propose R-ChFSI, a residual-based reformulation of ChFSI designed to exhibit strong convergence properties even when the matrix–vector products are computed inexactly. By recasting the Chebyshev polynomial recurrence in terms of residuals rather than direct eigenvector updates, R-ChFSI effectively suppresses errors introduced by approximate operator applications and preserves the convergence behaviour of the ChFSI scheme without approximations. We further derive convergence guarantees under matrix–vector product approximations, providing a rigorous foundation for the method in large-scale eigenvalue computations. The tolerance of R-ChFSI to inexact matrix-vector products enables an efficient treatment of generalized Hermitian definite eigenproblems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, where exact factorizations or high-accuracy iterative solves for evaluating $\textbf{B}^{-1}$ are often prohibitively expensive. By allowing the use of inexpensive approximate inverses of B, the method substantially reduces computational costs while maintaining robust convergence. Moreover, R-ChFSI naturally accommodates low-precision arithmetic for both standard and generalized eigenproblems, making it well-suited for modern hardware accelerators optimised for mixed-precision computation. To demonstrate the effectiveness of the approach, extensive numerical experiments are conducted on finite-element discretized eigenproblems with millions of degrees of freedom, solving for thousands of eigenpairs arising in ab initio material modelling using Kohn-Sham density functional theory. For generalized eigenproblems employing approximate $\textbf{B}^{-1}$, R-ChFSI achieves desired residual norms orders of magnitude smaller than those obtained with standard ChFSI. In addition, R-ChFSI reliably reaches target residual tolerances (e.g., 10-8) even with FP32 and TF32 arithmetic, significantly outperforming standard ChFSI in similar settings. By reducing dependence on high-precision arithmetic and enabling efficient use of modern accelerator hardware, R-ChFSI provides a scalable and computationally efficient eigensolver for large sparse eigenproblems central to ab initio electronic structure calculations, multiphysics simulations, and large-scale engineering analysis.
@@
Sparse Hermitian generalized eigenproblems, Chebyshev filtered subspace iteration, finite-element basis, Kohn-Sham density functional theory
journal: Nuclear Physics B
@@ 1 Introduction
Large sparse Hermitian eigenvalue problems play a central role in many areas of computational physics. They arise from the discretization of fundamental governing equations in quantum mechanics, electronic structure theory, fluid dynamics, plasma physics, wave propagation, electromagnetics, and elastodynamics. In such applications, the extremal eigenmodes often correspond to physically meaningful states—ground and low-lying excited states in quantum systems, dominant instability modes in fluid or plasma flows, or fundamental vibration frequencies in elastic media—which makes the efficient computation of only a small subset of the spectrum particularly important. The solution strategies for these eigenproblems often rely on fully iterative methods based on iterative orthogonal projection approaches. In these approaches, the large sparse matrix is orthogonally projected onto a carefully constructed smaller subspace rich in the wanted eigenvectors (Rayleigh-Ritz step), followed by subspace diagonalisation of the projected matrix and a subspace rotation step to recover the desired orthogonal eigenvector estimates of the original sparse Hermitian matrix. Popular iterative approaches include Davidson [crouzeixDavidsonMethod1994, davidsonIterativeCalculationFew1975], Generalized-Davidson [morganGeneralizationsDavidsonsMethod1986], Jacobi-Davidson [hochstenbachJacobiDavidsonMethod2006], Chebyshev-filtered subspace iteration (ChFSI) approach [zhouSelfconsistentfieldCalculationsUsing2006], LOBPCG [knyazevOptimalPreconditionedEigensolver2001], and PPCG [vecharynskiProjectedPreconditionedConjugate2015]. Another key class of iterative techniques is based on Krylov subspace methods, namely the Arnoldi method [arnoldiPrincipleMinimizedIterations1951], Lanczos methods [lanczosIterationMethodSolution1950] and their important variants, including implicit restart Arnoldi methods [sorensenImplicitApplicationPolynomial1992], Krylov-Schur method [stewartKrylovSchurAlgorithmLarge2002] and block-Krylov methods [saadNumericalMethodsLarge2011]. Our focus in this work involves the solution of large, sparse Hermitian eigenproblems using the ChFSI approach.
@@
Chebyshev filtered subspace iteration (ChFSI) has emerged as a robust alternative to Krylov subspace eigensolvers for extracting a small subset of extremal eigenpairs from large, sparse Hermitian matrices—particularly in scenarios where these eigenpairs must be computed repeatedly as the system matrix evolves within an outer iteration. A prominent example is the nonlinear eigenvalue problem that arises in Kohn-Sham density functional theory (DFT) [kohnSelfconsistentEquationsIncluding1965], where one is interested in the lowest eigenpairs of a Hamiltonian that depends nonlinearly on the eigenvectors. This problem is typically solved through a Self-Consistent Field (SCF) procedure, in which the Hamiltonian is updated at each iteration and a small set of extremal eigenpairs must be recomputed repeatedly as the system approaches self-consistency. ChFSI relies on a Chebyshev polynomial filtering procedure that constructs a subspace rich in the desired eigenvectors, exploiting the fast growth property of Chebyshev polynomials outside the interval [-1,1], followed by a Rayleigh-Ritz step of projection and subspace diagonalization. Approaches based on ChFSI have become quite popular recently to solve both standard and generalized eigenproblems arising in DFT [lohChebyshevFilteringLanczos1984, zhouParallelSelfconsistentfieldCalculations2006, motamarriHigherorderAdaptiveFiniteelement2013, levittParallelEigensolversPlanewave2015, wuAdvancingDistributedMultiGPU2023]. Although variants of ChFSI approaches combined with the Davidson method [zhouChebyshevDavidsonAlgorithm2007, miaoChebyshevDavidsonMethod2020, miaoFlexibleBlockChebyshevDavidson2023, wangNewSubspaceIteration2025] have been proposed, the ChFSI approach remains the preferred choice in electronic structure codes based on DFT for solving the underlying nonlinear eigenvalue problem [dasDFTFE10Massively2022, ghoshSPARCAccurateEfficient2017a, zhouChebyshevfilteredSubspaceIteration2014, fengMassivelyParallelImplementation2024, banerjeeChebyshevPolynomialFiltered2016] due to its scalability, ability to accommodate evolving subspaces and computational efficiency [dasFastScalableAccurate2019, dasLargeScaleMaterialsModeling2023]. Additionally, ChFSI is memory efficient compared to Davidson-type approaches and further offers an attractive alternative to preconditioned-type conjugate gradient approaches for problems, especially for problems where good preconditioners are unavailable. The current work fills an important gap in the ChFSI approach by developing a residual-based approach to ChFSI that is tolerant to inexact matrix-vector products during the subspace construction step. Consequently, we argue that the method is well-suited for solving generalized eigenproblems efficiently and additionally low-precision arithmetic can be leveraged in the light of recent changes in modern heterogeneous computing architectures, significantly improving computational efficiency.
@@
Modern hardware architectures have undergone substantial modifications in recent years due to the computational requirements of machine learning (ML) and artificial intelligence (AI) training. Owing to their high computational demands, these domains have gravitated towards the use of low-precision arithmetic for training and inferencing. In response to this demand, hardware manufacturers have been enhancing support for low-precision floating-point formats, such as tensorfloat32 and bfloat16, enabling significantly faster throughput and substantial performance improvements [^1] NVIDIA’s Blackwell GPUs, designed for AI/ML applications, demonstrate a notable decrease in peak double-precision (FP64) floating-point performance [HGXAISupercomputing] compared to the previous Hopper architecture.. These architectural changes underscore the necessity to modify scientific computing algorithms to efficiently utilize low-precision processes without compromising accuracy [dongarraHardwareTrendsImpacting2024, kashiMixedprecisionNumericsScientific2025]. For sparse eigensolvers, the ideas of mixed precision preconditioning for the LOBPCG eigensolver [kressnerMixedPrecisionLOBPCG2023] and mixed precision orthogonalization and Rayleigh-Ritz have been explored for the LOBPCG and ChFSI eigensolvers [kressnerMixedPrecisionLOBPCG2023, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020]. We note that in these works the evaluation of matrix-vector products involved in the subspace construction still need to be performed in the higher precision arithmetic. The development of robust iterative eigensolver algorithms to accommodate low precision arithmetic in matrix-vector multiplications becomes particularly important in light of the evolving hardware landscape.
@@
In this work, we introduce a residual-based reformulation of the Chebyshev filtered subspace iteration (ChFSI) method [saadChebyshevAccelerationTechniques1984, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, zhouChebyshevfilteredSubspaceIteration2014], referred to as the R-ChFSI method, for solving large-scale sparse Hermitian eigenvalue problems. The key novelty of R-ChFSI lies in its ability to accommodate inexact matrix-vector products in the subspace construction step while preserving convergence properties. Most importantly, for generalized eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, we demonstrate that the method naturally admits low-cost approximations of $\textbf{B}^{-1}$, allowing ChFSI to act on $\textbf{B}^{-1}\textbf{A}$ and solve large sparse generalized eigenproblems without degrading convergence, thereby avoiding expensive matrix factorizations and iterative solvers that are otherwise typically required for such problems [levittParallelEigensolversPlanewave2015]. The ability of R-ChFSI to be robust to approximations in matrix-vector multiplications makes the proposed method particularly relevant for modern hardware architectures that favor low-precision arithmetic, significantly improving computational efficiency. The core of our approach is a reformulated recurrence relation that modifies the standard ChFSI update step to operate on residuals rather than the guess of the eigenvectors. We then provide a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing inexact matrix-vector products compared to the traditional ChFSI recurrence relation. Our analysis establishes that the proposed R-ChFSI method effectively controls error propagation in the Chebyshev filtering process, resulting in more reliable convergence.
@@
Through extensive benchmarking on large-scale sparse generalized eigenvalue problems arising in ab initio calculations using finite-element (FE) discretization of DFT, where we employ a diagonal approximation to the FE basis overlap matrix inverse, R-ChFSI attains a residual tolerance that is orders of magnitude lower than standard ChFSI, demonstrating superior robustness to approximation errors. The improved convergence for generalized eigenvalue problems is particularly noteworthy, as it highlights the ability of R-ChFSI to accommodate approximate inverses while maintaining accuracy. In contrast, standard ChFSI is more sensitive to such approximations, often leading to a loss in accuracy or requiring significantly higher computational costs to achieve similar residual tolerances. This advantage of R-ChFSI is especially relevant in DFT calculations, where generalized eigenvalue problems frequently arise [levittParallelEigensolversPlanewave2015, ramakrishnanFastScalableFiniteelement2025, dasDFTFE10Massively2022], and high computational costs in the subspace construction constitute a significant bottleneck. Furthermore, R-ChFSI enables the use of lower-precision arithmetic in the filtering step, further enhancing the performance gains. In our benchmark studies, using TF32 arithmetic in Intel Data Center Max GPU accelerators deployed on the supercomuting system Aurora, we have obtained speedups of up to 2.3x for the filtering step and 1.9x for the complete eigensolve to reach the desired tolerance. Further, we have also employed reduced precision, BF16, for the nearest neighbor MPI communication arising in sparse matrix-vector multiplications to enhance the performance and we obtain speedups of up to 2.7x for the filtering step and 2.1x for the complete eigensolve to reach the desired tolerance. By reducing dependence on exact matrix factorizations and high-precision arithmetic, R-ChFSI enables efficient exploitation of modern hardware accelerators, making large-scale eigenvalue computations more feasible in high-performance computing environments.
@@
The remainder of this article is organized as follows: Section 2 outlines the key steps of the Chebyshev filtered subspace iteration procedure and subsequently analyzes the convergence properties of ChFSI in terms of how the maximum principal angle between the current subspace and the target eigenspace evolves during the iterations. Section 3 begins by analyzing the convergence of ChFSI when the subspace is constructed approximately due to inexact matrix-vector products. Subsequently, this section introduces the proposed residual-based Chebsyshev filtered subspace iteration method (R-ChFSI) for generalized eigenvalue problems. This section discusses the convergence of R-ChFSI with approximate matrix products and demonstrates mathematically that the R-ChFSI method can converge even when the traditional ChFSI method fails. Section 4 presents a comprehensive evaluation of the proposed R-ChFSI method in terms of accuracy and computational efficiency on GPU architectures that admit low precision arithmetic. It compares R-ChFSI to the traditional ChFSI approach for solving real symmetric and complex Hermitian eigenproblems in the case of generalized eigenproblems while using inexact matrix-vector products.
@@ 2 Mathematical Background
The ChFSI [saadChebyshevAccelerationTechniques1984, saadNumericalMethodsLarge2011, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, motamarriHigherorderAdaptiveFiniteelement2013, zhouChebyshevfilteredSubspaceIteration2014] approach for solving the desired eigenpairs belongs to the category of iterative orthogonal projection methods and is one of the widely used strategies to compute the smallest $n$ eigenvalues and their corresponding eigenvectors [dektorInexactSubspaceProjection2025, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020, zhangSPARCV200Spinorbit2024, ghoshSPARCAccurateEfficient2017a, xuSPARCSimulationPackage2021, kronikPARSECPseudopotentialAlgorithm2006, liouScalableImplementationPolynomial2020, winkelmannChASE2019, michaud-riouxRESCURealSpace2016, doganSolvingElectronicStructure2023, levittParallelEigensolversPlanewave2015]. To describe the proposed eigensolver strategy based on the ChFSI approach, in the current work, we consider the Hermitian generalized eigenvalue problem of the form

### v5

# Residual-based Chebyshev filtered subspace iteration for Hermitian eigenvalue problems tolerant to inexact matrix-vector products

@@ Abstract
Chebyshev Filtered Subspace Iteration (ChFSI) is widely used for computing a small subset of extremal eigenpairs from large matrices, particularly when the eigenpairs must be computed repeatedly as the system matrix evolves within an outer nonlinear iteration. These evolving subspaces are encountered, for instance, in Kohn-Sham density functional theory (DFT) used in quantum modelling of materials, subspace tracking in signal processing applications, and matrix completion problems. In this work, we propose R-ChFSI, a residual-based reformulation that recasts the Chebyshev polynomial recurrence in terms of residuals rather than eigenvector estimates, that achieves robust convergence even when matrix–vector products are computed inexactly. We derive convergence guarantees under such approximations and show that R-ChFSI can naturally leverage (i) the use of inexpensive approximate inverses for generalized eigenproblems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, where exact factorizations of B are prohibitively expensive, (ii) low-precision arithmetic (FP32, TF32) for both standard and generalized eigenproblems, and (iii) reduced-precision (BF16) inter-process communication in distributed sparse matrix–vector products. Controlled experiments on dense random matrices quantitatively verify the convergence bounds derived in this work and confirm the robustness of R-ChFSI to prescribed approximation errors for both standard and generalized eigenproblems. Large-scale experiments on finite-element discretized DFT generalized eigenproblems with up to 85 million grid points and 13,500 eigenpairs demonstrate that R-ChFSI achieves residual norms orders of magnitude below those of standard ChFSI when approximate inverses are employed, and reliably meets target tolerances of $10^{-8}$ even when employing reduced precision, yielding filtering speedups of up to $2.7{\times}$ ($2.1{\times}$ for the full eigensolver) on GPU accelerators.
@@
Hermitian generalized eigenproblems, Chebyshev filtered subspace iteration, finite-element basis, Kohn-Sham density functional theory
journal: Computer Physics Communications
@@ 1 Introduction
Large Hermitian eigenvalue problems play a central role in many areas of computational physics. They arise from the discretization of fundamental governing equations in quantum mechanics, electronic structure theory, fluid dynamics, plasma physics, wave propagation, electromagnetics, and elastodynamics. In such applications, the extremal eigenmodes often correspond to physically meaningful states—ground and low-lying excited states in quantum systems, dominant instability modes in fluid or plasma flows, or fundamental vibration frequencies in elastic media—which makes the efficient computation of only a small subset of the spectrum particularly important. The solution strategies for these eigenproblems often rely on fully iterative methods based on iterative orthogonal projection approaches. In these approaches, the large matrix is orthogonally projected onto a carefully constructed smaller subspace rich in the wanted eigenvectors (Rayleigh-Ritz step), followed by subspace diagonalisation of the projected matrix and a subspace rotation step to recover the desired orthogonal eigenvector estimates of the original Hermitian matrix. Popular iterative approaches include Davidson [5, 9], Generalized-Davidson [31], Jacobi-Davidson [17], Chebyshev-filtered subspace iteration (ChFSI) approach [53], LOBPCG [20], and PPCG [44]. Another key class of iterative techniques is based on Krylov subspace methods, namely the Arnoldi method [1], Lanczos methods [24] and their important variants, including implicit restart Arnoldi methods [40], Krylov-Schur method [41] and block-Krylov methods [38]. Our focus in this work involves the solution of large Hermitian eigenproblems using the ChFSI approach.
@@
Chebyshev filtered subspace iteration (ChFSI) has emerged as a robust alternative to Krylov subspace eigensolvers for extracting a small subset of extremal eigenpairs from large Hermitian matrices—particularly in scenarios where these eigenpairs must be computed repeatedly as the system matrix evolves within an outer iteration. A prominent example is the nonlinear eigenvalue problem that arises in Kohn-Sham density functional theory (DFT) [21], where one is interested in the lowest eigenpairs of a Hamiltonian that depends nonlinearly on the eigenvectors. This problem is typically solved through a Self-Consistent Field (SCF) procedure, in which the Hamiltonian is updated at each iteration and a small set of extremal eigenpairs must be recomputed repeatedly as the system approaches self-consistency. ChFSI relies on a Chebyshev polynomial filtering procedure that constructs a subspace rich in the desired eigenvectors, exploiting the fast growth property of Chebyshev polynomials outside the interval [-1,1], followed by a Rayleigh-Ritz step of projection and subspace diagonalization. Approaches based on ChFSI have become quite popular recently to solve both standard and generalized eigenproblems arising in DFT [27, 52, 32, 25, 48]. Although variants of ChFSI approaches combined with the Davidson method [54, 29, 28, 45] have been proposed, the ChFSI approach has become the preferred choice in large-scale electronic structure codes for solving the underlying nonlinear eigenvalue problem [8, 14, 51, 13, 2] due to its scalability, ability to accommodate evolving subspaces and computational efficiency [7, 6]. We note that planewave-based DFT codes often employ preconditioned iterative solvers (e.g., preconditioned conjugate gradient or Davidson-type methods) that exploit the availability of effective kinetic-energy-based preconditioners in the Fourier basis. However, for real-space discretizations (finite-element, finite-difference) where constructing such spectrally adapted preconditioners is less straightforward, ChFSI has proven to be the most effective approach for large-scale calculations. Additionally, ChFSI is memory efficient compared to Davidson-type approaches and further offers an attractive alternative to preconditioned-type conjugate gradient approaches for problems where good preconditioners are unavailable. The current work fills an important gap in the ChFSI approach by developing a residual-based approach to ChFSI that is tolerant to inexact matrix-vector products during the subspace construction step. Consequently, we argue that the method is well-suited for solving generalized eigenproblems efficiently and additionally low-precision arithmetic can be leveraged in the light of recent changes in modern heterogeneous computing architectures, significantly improving computational efficiency.
@@
Modern hardware architectures have undergone substantial modifications in recent years due to the computational requirements of machine learning (ML) and artificial intelligence (AI) training. Owing to their high computational demands, these domains have gravitated towards the use of low-precision arithmetic for training and inferencing. In response to this demand, hardware manufacturers have been enhancing support for low-precision floating-point formats, such as tensorfloat32 and bfloat16, enabling significantly faster throughput and substantial performance improvements [^1] NVIDIA’s Blackwell GPUs, designed for AI/ML applications, demonstrate a notable decrease in peak double-precision (FP64) floating-point performance [34] compared to the previous Hopper architecture.. These architectural changes underscore the necessity to modify scientific computing algorithms to efficiently utilize low-precision processes without compromising accuracy [12, 19]. For sparse eigensolvers, the ideas of mixed precision preconditioning for the LOBPCG eigensolver [22] and mixed precision orthogonalization and Rayleigh-Ritz have been explored for the LOBPCG and ChFSI eigensolvers [22, 8, 33]. We note that in these works the evaluation of matrix-vector products involved in the subspace construction still need to be performed in the higher precision arithmetic. The development of robust iterative eigensolver algorithms to accommodate low precision arithmetic in matrix-vector multiplications becomes particularly important in light of the evolving hardware landscape.
@@
In this work, we introduce a residual-based reformulation of the Chebyshev filtered subspace iteration (ChFSI) method [37, 52, 53, 51], referred to as the R-ChFSI method, for solving large-scale Hermitian eigenvalue problems. The key novelty of R-ChFSI lies in its ability to accommodate inexact matrix-vector products in the subspace construction step while preserving convergence properties. Most importantly, for generalized eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, we demonstrate that the method naturally admits low-cost approximations of $\textbf{B}^{-1}$, allowing ChFSI to act on $\textbf{B}^{-1}\textbf{A}$ and solve large generalized eigenproblems without degrading convergence, thereby avoiding expensive matrix factorizations and iterative solvers that are otherwise typically required for such problems [25]. The ability of R-ChFSI to be robust to approximations in matrix-vector multiplications makes the proposed method particularly relevant for modern hardware architectures that favor low-precision arithmetic, significantly improving computational efficiency. The core of our approach is a reformulated recurrence relation that modifies the standard ChFSI update step to operate on residuals rather than the guess of the eigenvectors. We then provide a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing inexact matrix-vector products compared to the traditional ChFSI recurrence relation. Our analysis establishes that the proposed R-ChFSI method effectively controls error propagation in the Chebyshev filtering process, resulting in more reliable convergence.
@@
Through extensive benchmarking on large-scale sparse generalized eigenvalue problems arising in ab initio calculations using finite-element (FE) discretization of DFT, where we employ a diagonal approximation to the FE basis overlap matrix inverse, R-ChFSI attains a residual tolerance that is orders of magnitude lower than standard ChFSI, demonstrating superior robustness to approximation errors. The improved convergence for generalized eigenvalue problems is particularly noteworthy, as it highlights the ability of R-ChFSI to accommodate approximate inverses while maintaining accuracy. In contrast, standard ChFSI is more sensitive to such approximations, often leading to a loss in accuracy or requiring significantly higher computational costs to achieve similar residual tolerances. This advantage of R-ChFSI is especially relevant in DFT calculations, where generalized eigenvalue problems frequently arise [25, 36, 8], and high computational costs in the subspace construction constitute a significant bottleneck. Furthermore, R-ChFSI enables the use of lower-precision arithmetic in the filtering step, further enhancing the performance gains. In our benchmark studies, using TF32 arithmetic in Intel Data Center GPU Max Series accelerators deployed on the Aurora supercomputing system, we have obtained speedups of up to 2.3x for the filtering step and 1.9x for the complete eigensolve to reach the desired tolerance. Further, we have also employed reduced precision, BF16, for the nearest-neighbour MPI communication arising in sparse matrix–vector multiplications to enhance performance, and we obtain speedups of up to 2.7x for the filtering step and 2.1x for the complete eigensolve to reach the desired tolerance. By reducing dependence on exact matrix factorizations and high-precision arithmetic, R-ChFSI enables efficient exploitation of modern hardware accelerators, making large-scale eigenvalue computations more feasible in high-performance computing environments.
@@
The remainder of this article is organized as follows: Section 2 outlines the key steps of the Chebyshev filtered subspace iteration procedure and subsequently analyzes the convergence properties of ChFSI in terms of how the maximum principal angle between the current subspace and the target eigenspace evolves during the iterations. Section 3 begins by analyzing the convergence of ChFSI when the subspace is constructed approximately due to inexact matrix-vector products. Subsequently, this section introduces the proposed residual-based Chebyshev filtered subspace iteration method (R-ChFSI) for generalized eigenvalue problems. This section discusses the convergence of R-ChFSI with approximate matrix products and demonstrates mathematically that the R-ChFSI method can converge even when the traditional ChFSI method fails. Section 4 presents a comprehensive evaluation of the proposed R-ChFSI method in terms of accuracy and computational efficiency on GPU architectures that admit low precision arithmetic. It compares R-ChFSI to the traditional ChFSI approach for solving real symmetric and complex Hermitian eigenproblems in the case of generalized eigenproblems while using inexact matrix-vector products.
@@ 2 Mathematical Background
The ChFSI [37, 38, 52, 53, 32, 51] approach for solving the desired eigenpairs belongs to the category of iterative orthogonal projection methods and is one of the widely used strategies to compute the smallest $n$ eigenvalues and their corresponding eigenvectors [10, 8, 33, 50, 14, 49, 23, 26, 47, 30, 11, 25]. To describe the proposed eigensolver strategy based on the ChFSI approach, in the current work, we consider the Hermitian generalized eigenvalue problem of the form